If a cup of ice cream is selected,what is the probability that the cholesterol content will be more than 670 mg?


1
Expert's answer
2021-06-03T07:20:38-0400

The average number of milligrams (mg) of cholesterol in a cup of certain brand of ice cream is 660 mg, and the standard deviation is 35 mg. assume the variable is normally distributed.

If a sample of 10 cups of ice cream is selected, what is the probability that the mean of the sample will be larger than 670 mg


Solution:

Given "\mu=660\ mg, \sigma=35\ mg, n=10"



"P(X>670)=1-P(X\leq670)""=1-P(Z\leq\dfrac{670-660}{35/\sqrt{10}})""\approx1-P(Z\leq0.9035)\approx0.1831"

The probability that the mean of the sample will be larger than 670 mg is 0.1831.